OpenStudy (anonymous):

Simplify the expression and show work please! (Come to the question, it can only be displayed as equation) Simplify the expression and show work please! (Come to the question, it can only be displayed as equation) @Mathematics

OpenStudy (anonymous):

\[\sqrt[3]{1/3}\]

OpenStudy (anonymous):

\[\sqrt[3]{1/6}\]

OpenStudy (anonymous):

Its the second one sorry.

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[3]{\frac{1}{6}}\] \[\Large \frac{\sqrt[3]{1}}{\sqrt[3]{6}}\] \[\Large \frac{1}{\sqrt[3]{6}}\] \[\Large \frac{1\times\sqrt[3]{6}\times\sqrt[3]{6}}{\sqrt[3]{6}\times\sqrt[3]{6}\times\sqrt[3]{6}}\] \[\Large \frac{\sqrt[3]{6\times6}}{\sqrt[3]{6\times6\times6}}\] \[\Large \frac{\sqrt[3]{36}}{\sqrt[3]{216}}\] \[\Large \frac{\sqrt[3]{36}}{6}\] So \[\Large \sqrt[3]{\frac{1}{6}}=\frac{\sqrt[3]{36}}{6}\]

OpenStudy (anonymous):

Thanks!!! You're my favorite math person! :P

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Can you help me with this one? \[3\div \sqrt[4]{144}\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{3}{\sqrt[4]{144}}\] \[\Large \frac{3}{\sqrt[4]{12^2}}\] \[\Large \frac{3}{(12^2)^{\frac{1}{4}}}\] \[\Large \frac{3}{12^{2\frac{1}{4}}}\] \[\Large \frac{3}{12^{\frac{2}{4}}}\] \[\Large \frac{3}{12^{\frac{1}{2}}}\] \[\Large \frac{3}{\sqrt{12}}\] \[\Large \frac{3\times\sqrt{12}}{\sqrt{12}\times\sqrt{12}}\] \[\Large \frac{3\sqrt{12}}{\sqrt{12\times12}}\] \[\Large \frac{3\sqrt{12}}{\sqrt{144}}\] \[\Large \frac{3\sqrt{12}}{\sqrt{12^2}}\] \[\Large \frac{3\sqrt{12}}{12}\] \[\Large \frac{3\sqrt{4\times3}}{12}\] \[\Large \frac{3\sqrt{4}\times\sqrt{3}}{12}\] \[\Large \frac{3\times2\times\sqrt{3}}{12}\] \[\Large \frac{6\sqrt{3}}{12}\] \[\Large \frac{\sqrt{3}}{2}\] So \[\Large \frac{3}{\sqrt[4]{144}} = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Thanks!!!!!!!!!!!! :D

jimthompson5910 (jim_thompson5910):

np